OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work y = x2 y = x + 2

OpenStudy (anonymous):

Set both right sides equal to each other: x^2 = x + 2 Now, put in standard form: x^2 - x - 2 = 0 now, factor: (x + 1)(x - 2) = 0 So, you have solutions at x = 2 and -1

OpenStudy (anonymous):

All good now, @topstryker ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Have a great day! You should, now that math is over!